OpenStudy (kc_kennylau):

This question is so fun :DDDDD What is the maximum \(k\) such that \(\large12^k|49!\)?

ganeshie8 (ganeshie8):

\(12^k | 49!\) \(\implies 49! = 12^k*m = (2^2\times 3)^k*m\)

OpenStudy (kc_kennylau):

I have to have dinner now, so I'll check you guys' answers when I finish my dinner :D

ganeshie8 (ganeshie8):

max k = least (#of 2^2 in 49!, #of 3 in 49!)

ganeshie8 (ganeshie8):

nice one :) i saw u working on similar question few days back... did u make this question ha ?

OpenStudy (kc_kennylau):

nope, I found it in http://brilliant.org/ but you didn't calculate the answer @ganeshie8

ganeshie8 (ganeshie8):

Okay, rest is trivial right #of 2^2 in 49! = (49/2 + 49/2^2 ... 59/2^5)/2 = (24 + 12 + 6 + 3 + 1)/2 = 23 #of 3 in 49! = (49/3 + 49/3^2 ... 59/2^3)/2 = (16 + 5+ 1)/2 = 22 max k = min(23, 22) = 22

ganeshie8 (ganeshie8):

u have a shortcut ? using wilson theorem or something ?

ganeshie8 (ganeshie8):

corrected typoes :) #of 2^2 in 49! = (49/2 + 49/2^2 ... 59/2^5)/2 = (24 + 12 + 6 + 3 + 1)/2 = 23 #of 3 in 49! = (49/3 + 49/3^2 ... 49/3^3) = (16 + 5+ 1) = 22 max k = min(23, 22) = 22

OpenStudy (kc_kennylau):

@ganeshie8 I'm back :D you forgot all the floor signs but other than that you're all correct :D

ganeshie8 (ganeshie8):

floor signs are implicit lol, dont be picky :P

OpenStudy (kc_kennylau):

lolz xD

OpenStudy (kc_kennylau):

and the plural of typo is typos

OpenStudy (kc_kennylau):

just fyi :D

ganeshie8 (ganeshie8):

c'mon

OpenStudy (kc_kennylau):

lolz sry :p

ganeshie8 (ganeshie8):

lol xD

ganeshie8 (ganeshie8):

wondering there must be a shorter way to knw the powers of a prime number in n!

OpenStudy (kc_kennylau):

don't think there's any :P

ganeshie8 (ganeshie8):

there must be, i feel lazy to do n/2, n/2^2.... humm

OpenStudy (kc_kennylau):

\[\Large\sum_{n=1}^{\left\lfloor\ln_2x\right\rfloor}\left\lfloor\frac{x}{2^n}\right\rfloor\]

OpenStudy (kc_kennylau):

this is the "shortcut" @ganeshie8

ganeshie8 (ganeshie8):

clever :|

OpenStudy (kc_kennylau):

lolz

OpenStudy (kc_kennylau):

or you can do so:

OpenStudy (kc_kennylau):